### v1

# A Mathematical Theory for Learning Semantic Languages by Abstract Learners

Abstract
Recent advances in Large Language Models (LLMs) have demonstrated the emergence of capabilities (learned skills) when the number of system parameters and the size of training data surpass certain thresholds. The exact mechanisms behind such phenomena are not fully understood and remain a topic of active research. Inspired by the skill-text bipartite graph model presented in [1] for modeling semantic language, we develop a mathematical theory to explain the emergence of learned skills, taking the learning (or training) process into account. Our approach models the learning process for skills in the skill-text bipartite graph as an iterative decoding process in Low-Density Parity Check (LDPC) codes and Irregular Repetition Slotted ALOHA (IRSA). Using density evolution analysis, we demonstrate the emergence of learned skills when the ratio of the size of training texts to the number of skills exceeds a certain threshold. Our analysis also yields a scaling law for testing errors relative to the size of training texts. Upon completion of the training, we propose a method for semantic compression and discuss its application in semantic communication.
I Introduction
In the recent era of natural language processing (NLP), the evolution of large language models (LLMs), such as GPT-4 [2] and Gemini [3], has greatly impacted people’s daily lives. There is a growing consensus that enhancing language model performance and sample efficiency across a broad spectrum of downstream NLP tasks is closely linked to scaling up these models. This scaling involves increasing both the size of the training data and the number of model parameters, as discussed in [4, 5].
The relationship between the scale of an LLM and its performance can often be quantitatively forecasted using scaling laws, a concept explored in-depth in [6, 7]. Despite this, not all downstream tasks exhibit continuous performance improvement with increased scale. This deviation from the expected trend, where certain tasks do not conform to predictive models, is highlighted in [8].
Further exploration in recent studies [9, 10, 2] demonstrates, through various numerical examples, that LLMs manifest emergent capabilities absent in their smaller counterparts. There is a notable enhancement in system performance once a certain critical scale threshold is surpassed. This phenomenon aligns with the concept of a phase transition as known in network science, a topic elaborately discussed in the book [11]. However, as pointed out in [9], the exact mechanisms behind the development of these emergent abilities in large language models are still not fully understood and remain a topic of active research.
In the study presented by [12], a simple explanation is offered for the phase transition phenomenon observed in LLMs based on the concept of list decoders. The approach in [12] models an LLM as a sequence-to-sequence random function over a certain token space with $M$ possible tokens. It was shown that the expected number of erroneous sequences in the list decoder can be bounded by a constant if $M\epsilon<1$, where $\epsilon$ is the false alarm probability. Since transformer-based LLMs with more parameters and extensive training can memorize more patterns [13, 14], they are more likely to reduce the false alarm probability $\epsilon$ below the percolation threshold of $1/M$. However, the list decoder approach does not explain how the capabilities of language skills emerge from scaling up the number of parameters and the size of training data.
In [1], a novel random graph approach was proposed to explain the scaling law discussed in [7]. A language is modeled by a random bipartite graph with skills on one side and texts (originally termed text-pieces in [1]) on the other side. An edge between a text and a skill indicates that understanding the text requires the skill. A text can be comprehended if all the skills connected to it are acquired.
Inspired by the random bipartite graph approach in [1], we take one step further by incorporating learning (or training). We summarize our contributions as follows:
(i) A mathematical theory for the emergence of learned skills: Specifically, we model the learning process for the skills in the skill-text bipartite graph as an iterative decoding process, similar to the analysis in Low-Density Parity Check (LDPC codes) [15, 16, 17] and Irregular Repetition Slotted ALOHA (IRSA) [18, 19, 20, 21, 22, 23]. When a text is presented to a learner, a skill in that text might be learned with a certain probability (specific definitions for learners are given in Section III). By repeatedly presenting a large number of training texts to a learner, a fraction of skills can be learned. Following the density evolution analysis for the iterative decoding process (in the asymptotic regime of a large number of skills) [24, 25, 26], we show the emergence of learned skills when the ratio $R$ of the number of training texts to the number of skills exceeds a certain threshold. Once this threshold is exceeded, the testing error, defined as the probability of whether the learner can understand a randomly selected text, drops sharply. Our analysis also provides the scaling law of the testing error with respect to the ratio $R$.
(ii) A semantic compression method: Two texts are said to be semantically equivalent if they require the same set of skills. A learner is considered generative if it can compose a semantically equivalent text for a set of learned skills. Once training for a generative learner is complete, the learner can be used as a semantic encoder/decoder. The idea is to index the learned skills and compress a text by using the indices of the skills required to understand that text. Since the testing error is very small when the threshold is exceeded, there is a very high probability that a randomly selected text can be compressed in this way. For texts not understood by the learner, we use traditional text compression methods.
The paper is organized as follows: In Section II, we define a semantic language using a skill-text bipartite graph. In Section III, we define abstract learners, including 1-skill learners in Section III-A and Poisson learners in Section III-B. Density evolution analysis is employed to derive the scaling law for these abstract learners. In Section IV, we demonstrate how the trained learners can be utilized for semantic compression and semantic communication. The paper concludes in Section V, where we discuss possible extensions of our work.
II Semantic languages
II-A Definition of a semantic language
In this section, we define the notion of semantic languages. This definition is inspired by the recent skill-text bipartite graph model presented in [1], which explains the emergence of new skills in language models.
Definition 1
(semantic language) A semantic language ${\cal L}=({\cal A},T,S,\phi)$ consists of (i) a set of tokens (symbols) ${\cal A}$, (ii) a set of texts $T$ composed of sequences of tokens, (iii) a set of skills $S$, and (iv) a function $\phi$ that maps a text $t$ in $T$ to a set of skills.
Two texts $t_{1}$ and $t_{2}$ are said to be semantically equivalent (or simply equivalent) if they both require the same set of skills, i.e., $\phi(t_{1})=\phi(t_{2})$. We say that a skill $s$ is present in a text $t$ if $s$ is included in the set of skills $\phi(t)$. For a semantic language ${\cal L}=({\cal A},T,S,\phi)$, we can generate a skill-text bipartite graph $G=(S,T,E)$, where an edge is added between a skill $s$ and a text $t$ if $s$ is present in $\phi(t)$ (see Figure 1 for an illustration). Note that the skill-text bipartite graph $G=(S,T,E)$ characterizes the semantic language.
Figure 1: An illustration for a skill-text bipartite graph, where the skill nodes connected to a text nodes indicates the set of skills required to understand the text. Two texts requiring the same set of skills are semnatically equivalent.
To motivate the definition of a semantic language, let us consider English as an example. The set of tokens in English comprises the characters that can be produced on a computer keyboard, typically encoded by the standard ASCII code. The set of English texts is vast, containing, as a subset, all the English texts available on the Internet. The set of skills required to understand these texts can also be very large and are generally latent and difficult to characterize.
In this paper, we address the question of learning a semantic language. A common approach is to sample a subset of texts from the set of texts $T$ and present this subset to a learner to learn the mapping $\phi$.
Definition 2
We define that a skill $s$ in a semantic language is learned if the learner can determine whether the skill $s$ is present in any given text $t$. We consider a text $t$ to be understood by a learner if all the skills contained within $t$ are learned by the learner.
II-B Learning a semantic language by sampling
As mentioned in the previous section, we sample a subset of $D$ texts from the set of texts $T$ and present these $D$ texts as the training texts. Let $R=D/|S|$ be the ratio of the number of training texts to the number of skills.
We make the following two assumptions about a randomly selected text:
(Poisson distribution) For a randomly selected text $t\in T$, the number of skills in $t$ follows a Poisson random variable with mean $c$, i.e., ${\bf\sf P}(|\phi(t)|=k)=e^{-c}\frac{c^{k}}{k!}$ for $k=0,1,2,\ldots$.
(Uniform distribution) The probability that a skill $s$ is included in a randomly selected text $t$ is uniformly distributed, i.e., ${\bf\sf P}(s\in\phi(t))=1/|S|$.
From (A1)-(A2), the skill-text bipartite graph of a sampled semantic language is a random bipartite graph with $|S|$ skill nodes on one side and $D$ text nodes on the other side. An edge exists between a skill node $s$ and a text node $t$ if $s\in\phi(t)$. The degree of a text node follows a Poisson distribution with mean $c$. According to (A1) and (A2), the number of edges between a text node and a particular skill node follows a Poisson distribution with mean $c/|S|$. Since there are $D$ text nodes, the number of edges of a skill node is also Poisson distributed, with mean $cD/|S|$. This leads to the following proposition.
Proposition 3
Consider the sampled skill-text bipartite graph from a semantic language ${\cal L}=({\cal A},T,S,\phi)$. Then: (i) The degree of a skill node follows a Poisson distribution with mean $cD/|S|=cR$. (ii) The probability that a randomly selected edge is connected to a particular text node is $1/D$, indicating a uniform distribution. Therefore, such a bipartite graph is a random bipartite graph, where the degree distribution of the text nodes (respectively, skill nodes) is Poisson with mean $c$ (respectively, $cR$).
III Abstract learners
In this section, we develop a mathematical theory for learning a semantic language using abstract learners.
To a learner, all skills are novel before training. After presenting a text $t$ to the learner, a skill $s$ in the text $t$ might be learned. After repeatedly presenting the training set of $D$ texts, we aim to determine the fraction of skills that are learned by the learner.
III-A 1-skill learners
Definition 4
(1-skill learner) An abstract learner is called a 1-skill learner if it can learn a novel skill $s$ by being presented with a text $t$ where the skill $s$ is the only novel skill in the text $t$. Once a skill $s$ is learned, the learner is able to identify the skill $s$ appearing in any other texts.
Now we describe how the 1-skill learner is trained. For each iteration, we present the $D$ sampled texts to the 1-skill learner in parallel. In the first iteration, a text containing only one skill is used to learn the skill in that text. Texts containing more than one skill do not contribute to learning in the first iteration. Once a skill $s$ is learned by the 1-skill learner, the number of novel skills in other texts containing $s$ is reduced by 1. In other words, the edges connected to the skill node $s$ can be removed from the skill-text bipartite graph. Therefore, the number of novel skills in the next iteration will not be greater than in the current iteration. In the second iteration, texts with only one novel skill are used to learn the skills in those texts. As in the first iteration, skills learned in the second iteration can be used to remove the corresponding edges in the skill-text bipartite graph. This iteration process is repeated until no more novel skills can be learned. We refer to this training process as Successive Cancellation of Novel Skills (SCNS).
This SCNS training process is mathematically equivalent to the iterative decoding approach in LDPC codes over the binary erasure channel (BEC) [15, 16, 17] and Irregular Repetition Slotted ALOHA (IRSA) [18, 19, 20, 21, 22, 23]. Following the density evolution analysis (in the asymptotic regime of a large number of skills) [24, 25, 26], let $q^{(i)}$ (respectively, $p^{(i)}$) be the probability that the skill end (respectively, the text end) of a randomly selected edge is not learned after the $i^{th}$ iteration of training. Initially, we have $q^{(0)}=1$. After the $i^{th}$ iteration, the number of novel skills in a randomly selected text follows a Poisson distribution with mean $q^{(i)}c$. For the 1-skill learner, a novel skill can be learned if it is the only novel skill in a text, which happens with probability $e^{-q^{(i)}c}$. Thus, we have
$p^{(i)}=1-e^{-q^{(i)}c}.$ (1)
Also, from SCNS, a skill end $s$ of a randomly selected edge is not learned after the $(i+1)^{th}$ iteration if the text nodes connected to $s$ (excluding the text end of the randomly selected edge) did not learn the skill $s$ after the $i^{th}$ iteration. Since the excess degree of a Poisson degree distribution is still Poisson with the same mean [11], the number of text nodes connected to $s$ (excluding the text end of the randomly selected edge) is a Poisson random variable with mean $cR$ (from Proposition 3(i)). Thus,
$q^{(i+1)}=\sum_{k=0}^{\infty}e^{-cR}\frac{(cR)^{k}}{k!}(1-p^{(i)})^{k}=e^{-cR(1-p^{(i)})}.$ (2)
Letting $i\to\infty$ in (1) and (2) yields
$p=1-e^{-ce^{-cR(1-p)}}.$ (3)
The equation in (3) characterizes the scaling law of the value $p$ with respect to $R$, the ratio of the number of training texts to the number of skills.
Now a skill $s$ remains novel after training if it cannot be learned from any text connected to $s$. As the degree of a skill node is Poisson distributed with mean $cR$, this occurs with probability $e^{-cR(1-p)}$. Thus, the probability that a randomly selected skill is learned is $p_{s}=1-e^{-cR(1-p)}$. For testing, we randomly select a text $t$ from $T$ (not included in the $D$ training texts). The number of skills in $t$ is also Poisson distributed with mean $c$. Then the probability that the text is understood by the learner is
$\sum_{k=0}^{\infty}e^{-c}\frac{c^{k}}{k!}(p_{s})^{k}=e^{-c(1-p_{s})}=e^{-ce^{-cR(1-p)}}.$ (4)
In view of (3), the testing error for the 1-skill learner is $1-p$, where $0\leq p\leq 1$ is the largest solution of (3). Since $p\geq 0$, we know from (4) that the testing error is at least $e^{-ce^{-cR}}$.
In Figure 2, we illustrate the effect of the parameters $R$ and $c$ on the values of $p$. We compute $p$ by using the recursive equation $p^{(i+1)}=1-e^{-(ce^{-cR(1-p^{(i)})})}$ with the initial condition $p^{(0)}=1$. The number of iterations is set to 10,000. Since $1-p$ also represents the probability that a (test) text is understood by the learner, there exists a percolation threshold, indicated by a sudden change in the values of $p$. For $c=3$ (respectively, 4 and 5), this threshold occurs when $R$ exceeds approximately 1 (respectively, 1.5 and 2.5). The underlying rationale is that for $c\geq 3$, a randomly chosen text is unlikely to encompass precisely one novel skill. Consequently, a substantial volume of training texts is essential to ensure a sufficient quantity of texts initially containing only a single novel skill. This foundation of learned skills subsequently initiates a cascade of successive cancellations of novel skills.
Figure 2: The plot illustrates the computed values of $p$ as a function of $R$, which ranges from 0.1 to 5 with a step size of 0.1, for each $c$ value (1, 2, 3, 4, 5). The computation is based on the recursive equation $p^{(i+1)}=1-e^{-(ce^{-cR(1-p^{(i)})})}$ with the initial condition $p^{(0)}=1$.
III-B Poisson learners
In this section, we extend the concept of the 1-skill learner to Poisson learners, where a novel skill can be learned with a certain probability.
Definition 5
(Poisson learner) An abstract learner is termed a Poisson learner if a novel skill $s$ in a text $t$ can be learned with probability $P_{\rm suc}(\rho)$ when the number of novel skills in that text is Poisson distributed with mean $\rho$. Once a skill $s$ is learned, the learner is capable of identifying the skill $s$ in any other texts.
The definition of Poisson learners is analogous to that of Poisson receivers in [27, 28, 29] for iterative decoders. As discussed in [27, 28, 29], a 1-skill learner is a Poisson learner with the success probability $P_{\rm suc}(\rho)=e^{-\rho}$. Similarly, the 2-skill learner, where skills in a text can be learned if the number of novel skills does not exceed two, is a Poisson learner with the success probability $P_{\rm suc}(\rho)=e^{-\rho}+\rho e^{-\rho}$ [27, 28, 29]. Using the density evolution method from [24, 25, 26], it was shown in [27, 28, 29] that
$p^{(i)}$ $=$ $1-P_{\rm suc}(q^{(i)}c),$ (5)
$q^{(i+1)}$ $=$ $e^{-cR(1-p^{(i)})}.$ (6)
Letting $i\to\infty$ yields
$p=1-P_{\rm suc}(ce^{-cR(1-p)}).$ (7)
Analogous to the argument for the 1-skill learner, the testing error for a Poisson learner is characterized in (4), where $0\leq p\leq 1$ is the largest solution in (7).
In Figure 3, we illustrate the effect of the parameters $R$ and $c$ on the values of $p$ for the 2-skill learner. We compute $p$ by using the recursive equation $p^{(i+1)}=1-P_{\rm suc}(ce^{-cR(1-p)})$ with the initial condition $p^{(0)}=1$, where $P_{\rm suc}(\rho)=e^{-\rho}+\rho e^{-\rho}$. The number of iterations is set at 10,000. A percolation threshold, indicated by a sudden change in the values of $p$, exists for $c=4$ (respectively, 5) when $R$ is approximately 0.5 (resp. 0.6). The percolation thresholds for the 2-skill learner are significantly lower than those for the 1-skill learner, demonstrating that the 2-skill learner is a much faster learner than the 1-skill learner.
Figure 3: The plot illustrates the computed values of $p$ as a function of $R$, which ranges from 0.1 to 5 with a step size of 0.1, for each $c$ value (1, 2, 3, 4, 5). The computation is based on the recursive equation $p^{(i+1)}=1-P_{\rm suc}(ce^{-cR(1-p)})$ with the initial condition $p^{(0)}=1$, where $P_{\rm suc}(\rho)=e^{-\rho}+\rho e^{-\rho}$.
We note that extending Poisson learners to settings with multiple classes of skills and texts is feasible. Such an extension is motivated by the diversity of subjects found in texts, including math, physics, chemistry, law, among others. The density evolution analysis for this extended scenario is akin to that for Poisson receivers with multiple classes of users and receivers, as detailed in [29]. A comprehensive analysis and numerical results are reserved for the journal version.
IV Semantic compression and semantic communication
Traditionally, lossless compression of a text requires exact recovery of the sequence of tokens in a text. However, if our interest lies in recovering only the semantic meaning of a text, we might be able to compress it using fewer bits than required for lossless compression. This concept is referred to as the semantic (level B) problem in [30].
Definition 6
(semantic compression) A compression method is termed a semantic compression method if the recovered text is semantically equivalent to the original text. An abstract learner is called generative if it can generate a text of tokens given a set of learned skills.
In this section, we assume that the abstract learners discussed in the previous section are also generative. Intuitively, to render an abstract learner generative, one can store a “table” of tokens corresponding to learned skills. Given a set of learned skills, the generative learner then uses the stored tokens to compose a text. It is known from [14] that transformers [13] are, in fact, Hopfield networks, which are a form of associative memory. In other words, transformers are capable of storing tables of tokens for learned skills.
Here, we demonstrate how to use generative learners for semantic compression. Once the training of a Poisson learner with the success probability function $P_{\rm suc}(\rho)$ is complete, the expected number of skills learned is $|S|(1-e^{-cR(1-p)})$, where $0\leq p\leq 1$ is the largest solution in (7). During the learning process, we can index the learned skills. Consequently, the number of bits required to represent a learned skill is $\log_{2}(|S|(1-e^{-cR(1-p)}))$. Therefore, if a text with $k$ skills is understood by the learner, it can be compressed/encoded using $k\log_{2}(|S|(1-e^{-cR(1-p)}))$ bits. As previously shown, the probability that a randomly selected text is understood by the learner is $e^{-(ce^{-cR(1-p)})}$, and it requires on average $c\log_{2}(|S|(1-e^{-cR(1-p)}))$ bits to encode the text. Conversely, if a randomly selected text is not understood by the learner, it can be encoded using a lossless compression encoder. Suppose the lossless compression encoder requires, on average, $z$ bits to compress a text. Then the semantic compression method described above requires, on average,
$e^{-(ce^{-cR(1-p)})}c\log_{2}(|S|(1-e^{-cR(1-p)}))+(1-e^{-(ce^{-cR(1-p)})})z$
bits for a text.
For semantic communication, we can utilize a semantic encoder/decoder. In Figure 4, we present the general architecture of a semantic communication system, which consists of a semantic transmitter, a semantic receiver, and a physical channel. The semantic transmitter includes a semantic encoder and a channel encoder, while the semantic receiver includes a semantic decoder and a channel decoder. The communication between the channel encoder and decoder occurs over the physical channel, referred to as level A communication in [30]. Conversely, the communication between the semantic encoder and decoder is not over the physical channel and is termed level B communication in [30].
Recent literature on semantic communications, such as [31, 32, 33], has proposed an end-to-end approach for jointly training the semantic and channel encoder/decoder. This method is claimed to be superior to separate training. However, the end-to-end approach does not scale efficiently with data size. As demonstrated in the previous section, a large dataset is necessary for training to exhibit the emergence of semantic capability. In light of this, a modular design akin to Internet protocol may be more effective for semantic communication.
Figure 4: The general framework of a semantic communication system.
V Conclusion
Inspired by [1], we defined a semantic language as a skill-text bipartite graph. Skills can be learned by repeatedly presenting training texts to abstract learners, such as 1-skill learners and Poisson learners. By using density evolution analysis, we showed the emergence of learned skills when the ratio $R$ exceeds a certain threshold. Moreover, we derived the corresponding scaling law for testing errors. We also demonstrated how trained abstract learners can be used for semantic compression and semantic communication.
However, there are still large gaps between our theory and real-world systems that require further study.
(i) Modeling abstract learners: Real-world LLMs, such as GPT-4 [2] and Gemini [3], are transformer-based and are trained using the gradient descent algorithm. It is not clear how a skill is learned when a text of tokens is fed into these transformers.
(ii) Connections to Kolmogorov complexity: In a recent talk [34], Sutskever pointed out the connection between compression and unsupervised learning through Kolmogorov complexity. It appears that GPT-4 has learned some structure of languages and uses that structure for semantic compression. However, it is not clear how such a structure is represented and learned.
(iii) Modeling the structure of a semantic language: In this paper, we use a skill-text bipartite graph to model the structure of a semantic language. This model is overly simplistic for real-world languages, which often have a hierarchical structure of skills. For example, basic skills must be learned before a complex skill can be mastered. Even though density evolution might be extended to settings with hierarchical skill structures, deriving simple scaling laws like those presented in this paper is unlikely.
(vi) Order of skills: if we view a skill as a thought, often a text can be represented by a chain (sequence) of thoughts. As such, the order of the skills appearing in a text could also be an important factor, thus, learning the mapping $\phi$ alone might not suffice for effective semantic encoding/decoding.
(v) Soft learning: in our model for an abstract learner, a skill is learned in a definite manner. Exploring a soft learning setting, where a skill is partially learned, could be beneficial. For transformer-based LLMs, a softmax function is used for the next token prediction. Extending our theory to encompass soft learning might help bridge the gap between the theory and the real-world systems.
(vi) Scaling laws of the number of parameters: The scaling law obtained in this paper only pertains to the size of training texts. It is unclear how testing errors scale with the number of parameters. For instance, GPT-3 is known to have more than 1,750 billion parameters, and LLMs prior to GPT-3 do not exhibit the emergence of skills. One possible explanation for the necessity of a larger number of parameters is that the capacity of a transformer, defined as the number of skills (or patterns) that can be stored in the corresponding Hopfield network [14], depends on the number of parameters in the transformer.
(vii) Experimental verification: performing experiments to verify our theory in the context of real-world semantic languages is likely impractical. However, it might be possible to construct a semantic language, e.g., a simple programming language, with a small number of skills. Subsequently, we could employ transformer-based LLMs with a relatively small number of parameters to verify the theory.