### v3

# A Mathematical Theory for Learning Semantic Languages by Abstract Learners

Abstract
Recent advances in Large Language Models (LLMs) have demonstrated the emergence of capabilities (learned skills) when the number of system parameters and the size of training data surpass certain thresholds. The exact mechanisms behind such phenomena are not fully understood and remain a topic of active research. Inspired by the skill-text bipartite graph model proposed by Arora and Goyal for modeling semantic languages, we develop a mathematical theory to explain the emergence of learned skills, taking the learning (or training) process into account. Our approach models the learning process for skills in the skill-text bipartite graph as an iterative decoding process in Low-Density Parity Check (LDPC) codes and Irregular Repetition Slotted ALOHA (IRSA). Using density evolution analysis, we demonstrate the emergence of learned skills when the ratio of the number of training texts to the number of skills exceeds a certain threshold. Our analysis also yields a scaling law for testing errors relative to this ratio. Upon completion of the training, the association of learned skills can also be acquired to form a skill association graph. We use site percolation analysis to derive the conditions for the existence of a giant component in the skill association graph. Our analysis can also be extended to the setting with a hierarchy of skills, where a fine-tuned model is built upon a foundation model. It is also applicable to the setting with multiple classes of skills and texts. As an important application, we propose a method for semantic compression and discuss its connections to semantic communication.
Index Terms:
Large language models, emergence of capabilities, Low-Density Parity Check codes, Irregular Repetition Slotted ALOHA, density evolution, semantic communication
I Introduction
In the recent era of natural language processing (NLP), the evolution of large language models (LLMs), such as GPT-4 [1] and Gemini [2], has greatly impacted people’s daily lives. There is a growing consensus that enhancing language model performance and sample efficiency across a broad spectrum of downstream NLP tasks is closely linked to scaling up these models. This scaling involves increasing both the size of the training data and the number of model parameters, as discussed in [3, 4]. The relationship between the scale of an LLM and its performance can often be quantitatively forecasted using scaling laws, as explored in [5, 6].
Recent studies [7, 8, 1] further demonstrates, through various numerical examples, that LLMs manifest emergent capabilities absent in their smaller counterparts. There is a notable enhancement in system performance once a certain critical scale threshold is surpassed, exhibiting a phase transition behavior often observed in network science [9]. However, as pointed out in [7], the exact mechanisms behind these emergent abilities in large language models are still not fully understood and remain a topic of active research.
In the study presented by [10], a simple explanation is offered for the phase transition phenomenon observed in LLMs based on the concept of list decoders. The approach in [10] models an LLM as a sequence-to-sequence random function over a certain token space with $M$ possible tokens. It was shown that the expected number of erroneous sequences in the list decoder can be bounded by a constant if $M\epsilon<1$, where $\epsilon$ is the false alarm probability. Since transformer-based LLMs with more parameters and extensive training can memorize more patterns [11, 12], they are more likely to reduce the false alarm probability $\epsilon$ below the percolation threshold of $1/M$. However, the list decoder approach does not explain how the capabilities of language skills emerge from scaling up the size of the training data. In [13], a novel random graph approach was proposed to explain the scaling law discussed in [6]. A language is modeled by a random bipartite graph with skills on one side and texts (originally termed text-pieces in [13]) on the other side. An edge between a text and a skill indicates that understanding the text requires the skill. A text can be comprehended if all the skills connected to it are acquired.
Inspired by the skill-text bipartite graph model proposed in [13], we develop a mathematical theory to explain the emergence of learned skills by an abstract learner during the learning (or training) process. We summarize our contributions as follows:
  • (i)
    Learning as an Iterative Decoding Process: We model the learning/training process for the skills in the skill-text bipartite graph as an iterative decoding process, similar to that adopted in Low-Density Parity Check (LDPC) codes [14, 15, 16] and Irregular Repetition Slotted ALOHA (IRSA) [17, 18, 19, 20, 21, 22]. When a text is presented to a learner, a skill in that text might be learned with a certain probability (specific definitions for learners are given in Section III). By repeatedly presenting a large number of training texts to a learner, a fraction of skills can be learned.
  • (ii)
    The Emergence of Learned Skills: Following the density evolution analysis for the iterative decoding process (in the asymptotic regime of a large number of skills) [23, 24, 25], we show the emergence of learned skills when the ratio $R$ of the number of training texts to the number of skills exceeds a certain threshold. Once this threshold is exceeded, the testing error, defined as the probability that a randomly selected text is not understood by the learner, drops sharply. Our analysis also provides the scaling law of the testing error with respect to the ratio $R$.
  • (iii)
    The Association of Skills: Upon completing training, a learner acquires not just isolated skills but also the interrelation of these skills, which can be represented by a skill association graph where edges connect learned skills appearing in the same text. Understanding this graph is crucial for making inferences, such as predicting the next skill from a given set. This process is akin to prompting a language model to forecast and generate a text based on a skill set. This leads to a critical inquiry: Are the learned skills sufficiently interconnected to aid in prediction? To address this question, we use site (or node) percolation analysis [9] to derive conditions for the existence of a giant component in the skill association graph.
  • (iv)
    Hierarchy of Skills: We consider a setting with two classes of skills: basic and domain-specific. Learning a domain-specific skill first requires acquiring a random number of basic skills, establishing a hierarchy of skills. Initially, we train a foundation model on basic skills, which is then fine-tuned with domain-specific texts. Our findings identify two critical thresholds for minimizing testing errors in domain-specific tasks: (i) a large number of basic skills are learned when the number of basic texts exceeds the threshold in the foundation model, and (ii) a large number of domain-specific skills are learned when the number of domain-specific texts exceeds the threshold in the fine-tuning model.
  • (v)
    Multiple Classes of Skills and Texts: We extend the density evolution analysis to the settings with multiple classes of skills and texts, driven by the diversity of subjects in texts like math, physics, chemistry, law, etc. Such an extension is rooted on the framework of Poisson receivers with multiple classes of users and receivers in [26]. Interestingly, thresholds are also observed for the setting with multiple classes of skills and texts. The thresholds for various skill classes align due to the coupling of the density evolution equations, as illustrated in [26], resulting in the simultaneous emergence of various classes of skills.
  • (vi)
    Application to Semantic Compression: We propose a semantic compression scheme that encodes the learned skills instead of all words in a text. In particular, we define the notion of a generative learner that is trained to identify the set of skills in a semantic language. Once trained, it can be used to perform semantic compression by representing a text using indices of the learned skills required to understand that text.
The paper is organized as follows. In Section II, we define a semantic language using a skill-text bipartite graph. In Section III, we introduce abstract learners, including 1-skill learners in Section III-A and Poisson learners in Section III-B. Density evolution analysis is employed to derive the scaling law for these abstract learners. In Section IV, we address the problem of learning the association between two learned skills and derive conditions for the existence of a giant component in the skill association graph. In Section V, we further address the problem setting with a hierarchy of skills by using a foundation model and a fine-tuning model. We present the extension to the setting with multiple classes of skills and texts in Section VI. In Section VII, we demonstrate how the trained learners can be utilized for semantic compression and semantic communication. The paper concludes in Section VIII, where we discuss possible extensions of our work.
II Semantic Languages
II-A Definition of a Semantic Language
In this section, we define the notion of semantic languages. This definition is inspired by the recent skill-text bipartite graph model presented in [13], which explains the emergence of new skills in language models.
Definition 1
[…]
To motivate the definition of a semantic language, let us consider English as an example. The set of tokens in English comprises the characters that can be produced on a computer keyboard, typically encoded by the standard ASCII code. The set of English texts is vast, containing, as a subset, all the English texts available on the Internet. The set of skills required to understand these texts can also be very large and are generally latent and difficult to characterize.
In this paper, we address the question of learning a semantic language. A common approach is to sample a subset of texts from the set of texts ${\cal T}$ and present this subset to a learner to learn the mapping $\phi$.
Definition 2
A skill $s$ in a semantic language is learned if the learner can determine whether the skill $s$ is present in any given text $t$ in $𝒯\color[rgb]{0,0,0}{\cal T}$. A text $t$ is understood by a learner if all the skills contained within $t$ are learned by the learner.
II-B Learning a Semantic Language by Sampling
Suppose that a subset of texts ${\cal D}$ is sampled from the set of texts $𝒯\color[rgb]{0,0,0}{\cal T}$ and presented to the learner as training texts. Let $𝒮R=|{\cal D}|/|\color[rgb]{0,0,0}{\cal S}|$ be the ratio of the number of training texts to the number of skills.
We make the following two assumptions about the sampled texts:
(Poisson degree distribution) Let $N(t)=|\phi(t)|$ be the number of skills in the text $t\in{\cal D}$. Then $\{N(t),t\in{\cal D}\}$ are independent Poisson random variables with mean $c$, i.e., ${\bf\sf P}(N(t)=k)=e^{-c}\frac{c^{k}}{k!}$ for $k=0,1,2,\ldots$.
(Uniform connection) Each edge of a text node is connected to a skill node independently and uniformly. Thus, the probability that an edge of the text node $t$ is connected to a particular skill node $s$ is $1/|{\cal S}|$.
Figure 1: An illustration for a skill-text bipartite graph, where the skill nodes connected to a text node indicates the set of skills required to understand the text. Two texts requiring the same set of skills are semantically equivalent.
From (A1) and (A2), the skill-text bipartite graph of a sampled semantic language is a random bipartite graph with $|{\cal S}|$ skill nodes on one side and $|{\cal D}|$ text nodes on the other side. Denote by $M(s,t)$ the number of edges between a skill node $s$ and a text node $t$. Since randomly splitting a Poisson process yields independent Poisson processes, we have from (A1) and (A2) that $\{M(s,t),s\in{\cal S},t\in{\cal D}\}$ are independent Poisson random variables with mean $c/|{\cal S}|$. Denote by $M(s)$ the number of edges of a skill node $s$. As a superposition of independent Poisson processes (from the $|{\cal D}|$ text nodes) is a Poisson process, we know that $\{M(s),s\in{\cal S}\}$ are independent Poisson random variables with mean $c|{\cal D}|/|{\cal S}|$ (i.e., $cR$). On the other hand, the probability that a point in a superposition of independent Poisson processes comes from one of these processes is proportional to their rates. This leads to the following proposition.
Proposition 3
Consider a sampled skill-text bipartite graph from a semantic language ${\cal L}=({\cal A},{{\cal T},{\cal S},}\phi)$. Then:
(Poisson degree distribution) $\{M(s),s\in{\cal S}\}$ are independent Poisson random variables with mean $c|{\cal D}|/|{\cal S}|=cR$, where $M(s)$ is the degree of the skill node $s$.
(Uniform connection) Each edge of a skill node is connected to a text node independently and uniformly. Thus, the probability that an edge of a skill node is connected to a particular text node is $1/|{\cal D}|$.
Therefore, such a bipartite graph is a random bipartite graph, as per the configuration model in [9], where the degree distribution of the text nodes (respectively, skill nodes) is Poisson with mean $c$ (respectively, $cR$).
III Abstract Learners
In this section, we develop a mathematical theory to understand the learning process of a semantic language by an abstract learner. To a learner, all skills are novel before training, but a skill $s$ in text $t$ may be learned after the text is presented to the learner. By repeatedly presenting a set of training texts ${\cal D}$, we aim to determine the fraction of skills that can be learned by the learner. We consider two types of learners based on their learning capabilities, namely, a 1-skill learner and a more general Poisson learner.
III-A 1-Skill Learner
Definition 4
(1-Skill Learner) An abstract learner is called a 1-skill learner if it can learn a novel skill $s$ by being presented with a text $t$ where the skill $s$ is the only novel skill in the text. Once a skill $s$ is learned, the learner is able to identify the skill $s$ appearing in any other text.
Now we describe how the 1-skill learner is trained. For each iteration, we present the $|{\cal D}|$ sampled texts to the 1-skill learner in parallel. In the first iteration, a text containing only one skill is used to learn the skill in that text. Texts containing more than one skill do not contribute to learning in the first iteration. Once a skill $s$ is learned by the 1-skill learner, the number of novel skills in other texts containing $s$ is reduced by 1. In other words, the edges connected to the skill node $s$ can be removed from the skill-text bipartite graph. Therefore, the number of novel skills associated with each text in the next iteration will not be greater than that in the current iteration. In the second iteration, texts with only one novel skill remaining are used to learn the skills in those texts. As in the first iteration, skills learned in the second iteration can be used to remove the corresponding edges in the skill-text bipartite graph. This iteration process is repeated until no more novel skills can be learned. We refer to this training process as the Successive Cancellation of Novel Skills (SCNS).
Figure 2: An example of the Successive Cancellation of Novel Skills (SCNS).
In Figure 2, we present a simple example of SCNS. Figure 2(a) depicts the original skill-text bipartite graph. Figure 2(b) focuses on the skill node $a$, which is connected to a text node with degree $1$; hence, we use a red edge to indicate that the skill node $a$ is learned by the text node. Consequently, the skill node $a$ can be learned, allowing the edges connected to it to be removed from the skill-text bipartite graph. Figure 2(c) shows the graph after the edges are removed, revealing that the skill node $b$ is connected to a text node with degree $1$, indicating it can also be learned. Figure 2(d) represents the last iteration, where no additional skills can be learned, concluding the process with the skills $a$ and $b$ being learned, while skills $c$ and $d$ are not.
This SCNS training process is mathematically equivalent to the iterative decoding approach in LDPC codes over the binary erasure channel (BEC) [14, 15, 16] and Irregular Repetition Slotted ALOHA (IRSA) [17, 18, 19, 20, 21, 22]. Following the density evolution analysis (in the asymptotic regime of a large number of skills) [23, 24, 25], we can derive the probability that a randomly selected text is not understood by the learner. This can be viewed as the training error of the learner as a result of the above training process. In particular, let $q^{(i)}$ be the probability that the skill end of a randomly selected edge is not learned after the $i^{th}$ SCNS iteration of training. Initially, we have $q^{(0)}=1$ (as all the skills are novel before training). After the $i^{th}$ SCNS iteration, the number of novel skills in a randomly selected text follows a Poisson distribution with mean $q^{(i)}c$ (a more detailed argument for this is provided in Section VI-A). For the 1-skill learner, a novel skill can be learned if it is the only novel skill in a text, which happens with probability $e^{-q^{(i)}c}$. Thus, the probability that the text end of a randomly selected edge does not contribute to the learning of a novel skill in the $(i+1)^{th}$ SCNS iteration is
$p^{(i+1)}=1-e^{-q^{(i)}c}.$ (1)
To compute $q^{(i+1)}$, let us consider a randomly selected edge. Suppose that the skill (resp. text) end of that edge is $s$ (resp. $t$). By SCNS, the skill $s$ is not learned after the $(i+1)^{th}$ SCNS iteration if the text nodes connected to $s$ (excluding the text node $t$ from the randomly selected edge) are not able to contribute to the learning of skill $s$ in the $(i+1)^{th}$ SCNS iteration. Since the excess degree of a Poisson degree distribution is still Poisson with the same mean [9], the number of text nodes connected to $s$ (excluding the text node $t$) is a Poisson random variable with mean $cR$ (from Proposition 3(i)). Thus,
$q^{(i+1)}=\sum_{k=0}^{\infty}e^{-cR}\frac{(cR)^{k}}{k!}{p^{(i+1)}}^{k}=e^{-cR(1-p^{(i+1)})}.$ (2)
Letting $i\to\infty$ in (1) and (2) yields
$p=1-e^{-ce^{-cR(1-p)}}.$ (3)
Here, $p$ represents the probability that the text end of a randomly selected edge still contains more than one novel skill at the end of the training process (i.e., after the learner has been repeatedly presented with the set of training texts ${\cal D}$). Equation (3) characterizes the scaling law of the value $p$ with respect to $R$, the ratio of the number of training texts to the number of skills.
Now a skill $s$ remains novel after training if it cannot be learned from any text connected to $s$. (see Figure 3 for an illustration). As the degree of a skill node is Poisson distributed with mean $cR$, this occurs with probability $e^{-cR(1-p)}$. Thus, the probability that a randomly selected skill is learned is
$\zeta=1-e^{-cR(1-p)}.$ (4)
Figure 3: We illustrate the density evolution (tree) analysis by considering a skill node $s$. The skill node $s$ remains novel if it cannot be learned from any text connected to $s$. This leads to (4) (as the degree of node $s$ is Poisson with mean $cR$). For the 1-skill learner, the text end $t$ of an edge connected to $s$ can learn $s$ if it is the only edge connected to $t$, i.e., all the other edges connected to $t$ (excluding the edge between $s$ and $t$) are removed. This leads to (1) (as the excess degree of node $t$, excluding the edge between $s$ and $t$, is Poisson with mean $c$and each excess edge is removed with probability $q$). Now let us go down the tree one level further. The edge between $t$ and $s^{\prime}$ cannot be removed if $s^{\prime}$ is not learned (remains novel). The skill node $s^{\prime}$ remains novel if it cannot be learned from any text connected to $s^{\prime}$ (excluding the text $t$). This leads to (2) (as the excess degree of node $s^{\prime}$, excluding the edge between $s^{\prime}$ and $t$, is Poisson with mean $cR$).
For testing, we randomly select a text $t$ from ${\cal T}$ (not included in the set of training texts ${\cal D}$). Assume that this randomly selected text has the same distribution as those in the training data, i.e., it also satisfies (A1) and (A2). As such, the number of skills in $t$ is also Poisson distributed with mean $c$. Recall that a text can be understood if all the connected skill nodes are learned (see Figure 4 for an illustration). Define the testing error (denoted by $\epsilon$) as the probability that the randomly selected text is not understood by the learner. For the 1-skill learner, we have
$\epsilon$ $=$ $1-\sum_{k=0}^{\infty}e^{-c}\frac{c^{k}}{k!}\zeta^{k}$ (5)
$=$ $1-e^{-c(1-\zeta)}=1-e^{-ce^{-cR(1-p)}}.$
In this case, the probability of the testing error is also equal to the solution for $p$ in (3), for $0\leq p\leq 1$. Since $p\geq 0$, we know from (5) that the testing error is at least $1-e^{-ce^{-cR}}$.
Figure 4: A randomly selected text can be understood if all the connected skill nodes are learned (highlighted in red circles). This leads to (5).
Figure 5: The plot illustrates the probability of the testing error $\epsilon$ as a function of $R$ for the 1-skill learner.
In Figure 5, we illustrate the effect of the parameters $R$ and $c$ on the probability of the testing error $\epsilon$. We compute $p$ by using the recursive equation $p^{(i+1)}=1-e^{-ce^{-cR(1-p^{(i)})}}$ with the initial condition $p^{(0)}=1$. The number of iterations is set to 10,000. We can see that there exists a percolation threshold, indicated by a sudden change in the values of $\epsilon$. The phase transition behavior is most evident when the average number of skills contained in each text (i.e., $c$) is greater than $3$. For $c=3$ (respectively, 4 and 5), this threshold occurs when $R$ exceeds approximately 1 (respectively, 1.5 and 2.5). The underlying rationale is that for $c\geq 3$, a randomly chosen text is unlikely to encompass precisely one novel skill. Consequently, a substantial volume of training texts is essential to ensure a sufficient quantity of texts initially containing only a single novel skill. This foundation of learned skills subsequently initiates a cascade of successive cancellations of novel skills.
III-B Poisson Learner
In this section, we extend the concept of the 1-skill learner to the Poisson learner, where multiple novel skills may be learned from a text with a certain probability.
Definition 5
(Poisson Learner) An abstract learner is termed a Poisson learner if a novel skill $s$ in a text $t$ can be learned with probability $P_{\rm suc}(\rho)$ when the number of novel skills in that text is Poisson distributed with mean $\rho$. Once a skill $s$ is learned, the learner is able to identify the skill $s$ appearing in any other text.
The definition of Poisson learners is analogous to that of Poisson receivers in [27, 28, 26] for iterative decoders. As discussed in [27, 28, 26], a 1-skill learner is a Poisson learner with the success probability $P_{\rm suc}(\rho)=e^{-\rho}$. Similarly, the 2-skill learner, where skills in a text can be learned if the number of novel skills does not exceed two, is a Poisson learner with the success probability $P_{\rm suc}(\rho)=e^{-\rho}+\rho e^{-\rho}$ [27, 28, 26]. Using the density evolution method from [23, 24, 25], it was shown in [27, 28, 26] that
$p^{(i+1)}$ $=$ $1-P_{\rm suc}(q^{(i)}c),$ (6)
$q^{(i+1)}$ $=$ $e^{-cR(1-p^{(i+1)})}.$ (7)
Letting $i\to\infty$ yields
$p=1-P_{\rm suc}(ce^{-cR(1-p)}).$ (8)
Then, analogous to the argument for the 1-skill learner, the testing error for a Poisson learner is characterized in (5) with $p$ given in (8).
Figure 6: The plot illustrates the probability of the testing error $\epsilon$ as a function of $R$ for the 2-skill learner.
In Figure 6, we illustrate the effect of the parameters $R$ and $c$ on the probability of the testing error $\epsilon$ for the 2-skill learner. We first compute $p$ by using the recursive equation $p^{(i+1)}=1-P_{\rm suc}(ce^{-cR(1-{p^{(i)}})})$ with the initial condition $p^{(0)}=1$ and $P_{\rm suc}(\rho)=e^{-\rho}+\rho e^{-\rho}$. The number of iterations is set at 10,000. A percolation threshold, indicated by a sudden change in the values of $\epsilon$, exists for $c=4$ (respectively, 5) when $R$ is approximately 0.5 (resp. 0.6). The percolation thresholds for the 2-skill learner are significantly lower than those for the 1-skill learner, demonstrating that the 2-skill learner is a much faster learner than the 1-skill learner.
IV Learning the association of skills
When the training is complete, a learner can learn not only a fraction of skills but also the associations between these learned skills. We say two learned skills are associated if they appear in the same text. Since a learned skill can be identified whenever it is present in a text, we can generate a skill association graph for learned skills by adding an edge between two learned skill nodes if they appear in the same text. For instance, the two learned skills, $a$ and $b$, shown in Figure 2, are associated as they appear in the same text. After training, we can generate a skill association graph with nodes $a$ and $b$, and an edge connecting these two nodes.
Knowing the structure of the skill association graph is crucial, as it can be utilized for inference purposes, such as predicting the next skill given a set of skills. This is analogous to giving an LLM a prompt in the form of a text with a set of skills and asking the LLM to predict the next skill and generate a text based on the predicted skill. An essential question arises: Are most of the learned skills interconnected in a way that facilitates their use for prediction? This question guides us to perform a giant component analysis of the skill association graph. Note that a giant component in a random graph is a connected subgraph whose size is proportional to the size of the graph. It is well-known [9] that there exists at most one giant component in a random graph. The rest of the components are called small components. One important property of a random graph is that small components are trees with high probability.
Our approach to this problem is based on the site (or node) percolation analysis as described in [9]. As in Section III-B, we consider the skill-text bipartite graph with the set of skill nodes ${\cal S}$ and the set of text nodes ${\cal D}$. It is important to note that not every skill is learned upon the completion of training. In fact, as shown in (4), the probability that a randomly selected skill node is learned is given by
$\zeta=1-e^{-cR(1-p)},$ (9)
where $p$ is the solution of (8).
To apply the site percolation analysis for the skill-text bipartite graph, we assume that skill nodes are learned independently with probability $\zeta$. Let $\mu_{s}$ be the probability that a skill node is connected to a small component via one of its edges. Also, let $\mu_{t}$ be the probability that a text node is connected to a small component via one of its edges. Recall from Proposition 3 that the degree distribution of the text nodes (respectively, skill nodes) is Poisson with mean $c$ (respectively, $cR$). Note that a skill node is only connected to a small component via one of its edges (called the tagged edge) if all the excess edges of the text end of the tagged edge are only connected to small components. Since the excess degree distribution of a Poisson degree distribution is also Poisson with the same mean, we then have (from the tree property for small components) that
$\mu_{s}=\sum_{k=0}^{\infty}e^{-c}\frac{c^{k}}{k!}\mu_{t}^{k}=e^{-c(1-\mu_{t})}.$ (10)
Now we derive the governing equation for $\mu_{t}$. If the skill end of an edge (called the tagged edge) is not learned, then there are no further connected skill nodes via this edge, which occurs with probability $1-\zeta$. Conversely, if the skill end of the tagged edge is learned, then it is only connected to a small component if all the excess edges of the skill end of the tagged edge are only connected to small components. Thus, we have (from the tree property for small components) that
$\mu_{t}$ $=$ $(1-\zeta)+\zeta\sum_{k=0}^{\infty}e^{-cR}\frac{{(cR)}^{k}}{k!}\mu_{s}^{k}$ (11)
$=$ $(1-\zeta)+\zeta e^{-Rc(1-\mu_{s})}.$
Clearly, $\mu_{s}=\mu_{t}=1$ is a trivial solution of (10) and (11). To find a nontrivial solution (if any), one can iterate the following two equations:
$\mu_{s}^{(i+1)}=e^{-c(1-\mu_{t}^{(i)})},$
$\mu_{t}^{(i+1)}=(1-\zeta)+\zeta e^{-Rc(1-\mu_{s}^{(i)})},$ (12)
with the initial condition $\mu_{s}^{(0)}=\mu_{t}^{(0)}=0$. By induction, it is easy to see from (IV) that $0\leq\mu_{s}^{(i)}<\mu_{s}^{(i+1)}\leq 1$ and $0\leq\mu_{t}^{(i)}<\mu_{t}^{(i+1)}\leq 1$ for all $i$. Thus, the recursion in (IV) converges to a solution in $[0,1]\times[0,1]$ as $i\to\infty$.
Note that a randomly selected skill node is in the giant component if the skill node is learned and at least one of its edges is not connected to a small component. Thus, the probability that a randomly selected skill node is in the giant component, denote by $p_{G}$, is
$p_{G}={\zeta}\Big{(}1-\sum_{k=0}^{\infty}e^{-cR}\frac{{(cR)}^{k}}{k!}\mu_{s}^{k}\Big{)}=\zeta(1-e^{-Rc(1-\mu_{s})}).$ (13)
Figure 7: The plot of the probability $p_{G}$ that a randomly selected skill is in the giant component for the 1-skill learner as a function of $R$ for various values of $c$.
In Figure 7, we present a plot of the probability $p_{G}$ that a randomly selected skill is in the giant component as a function of $R$ for the 1-skill learner with values of $c=1,2,3,4,5$. The solid line represents the theoretical value, and the dashed line represents the simulated value obtained by averaging over 100 random bipartite graphs with a total of 20,000 skill and text nodes. This figure offers an intriguing comparison to the testing errors shown in Figure 5. For $c=1$, as indicated in Figure 5, the testing errors decrease with an increase in $R$. However, the learned skills do not form a giant component until $R$ surpasses a threshold, approximately at 1.5, as depicted in Figure 7. While it is relatively easy for the 1-skill learner to acquire skills when the average number of skills in a sampled text is small, it is also challenging to learn the association between two learned skills. Conversely, for $c=5$, learning individual skills is difficult. Yet, when $R$ exceeds a threshold, roughly at 2.5, we observe a significant reduction in testing errors in Figure 5. Additionally, surpassing this same threshold in Figure 7 correlates with the emergence of a giant component with the size near the size of the skill nodes. This suggests that the acquisition of skills and their associations occur simultaneously once this threshold is exceeded. It is noteworthy that discrepancies exist between the theoretical and simulation results, as seen in Figure 7 for $c=1$ and $2$. These discrepancies arise because the percolation analysis assumes skill nodes are learned independently–a premise that does not hold when $c$ is small.
Next, we derive the condition for the existence of a giant component. Combining (10) and (11) yields
$\mu_{s}$ $=$ $e^{-c\{1-[(1-\zeta)+\zeta e^{-Rc(1-\mu_{s})}]\}}$ (14)
$=$ $e^{-c\zeta(1-e^{-Rc(1-\mu_{s})})}.$
Consider the function
$g_{1}(\mu_{s})=e^{-c\zeta(1-e^{-Rc(1-\mu_{s})})}.$ (15)
Note that $g_{1}(\mu_{s})$ is a convex function of $\mu_{s}$ when $0\leq\mu_{s}\leq 1$. This can be verified by showing that the second derivative of $g_{1}(\mu_{s})$ is nonnegative for $0\leq\mu_{s}\leq 1$. Clearly, $\mu_{s}=1$ is the trivial solution of (14). By the convexity and the fact that $g_{1}(0)>0$, we have a non-trivial solution for some $\mu_{s}<1$ if the slope of $g_{1}(\mu_{s})$ at $\mu_{s}=1$ is greater than the slope of the line $y=\mu_{s}$, i.e.,
$\left.{\frac{d}{d\mu_{s}}g_{1}(\mu_{s})}\right|_{\mu_{s}=1}>1.$ (16)
This condition is equivalent to
$c^{2}R\zeta>1$ (17)
or, by the representation of $\zeta$ in (9),
$c^{2}R(1-e^{-cR(1-p)})>1.$ (18)
Here, $p$ is the probability that the text end of a randomly selected edge is not learned and is given by the solutions of (3) (respectively, (8)) for a 1-skill learner (respectively, Poisson learner). In Figure 8, we provide an illustration of the condition in (16) for $c=3$ and $R=0.5,1,1.5$ by considering a 1-skill learner. We can see that a non-trivial solution exists at the intersection between the curve $y=g_{1}(\mu_{s})$ and the line $y=\mu_{s}$ for all these cases.
Figure 8: The solution of the equation (14) is found at the point where the curve $y=g_{1}(\mu_{s})$ intercepts the line $y=\mu_{s}$, given that $c=3$.
Figure 9: Regions based on the slope of the function $g_{1}(\mu_{s})$ at $\mu_{s}=1$. The red region corresponds to $g_{1}(\mu_{s})>1$, indicating the existence of a giant component, and the blue region corresponds to $g_{1}(\mu_{s})<1$, signifying the absence of a giant component.
For each pair of $c$ and $R$, we can check whether the condition in (18) is satisfied. If it is satisfied, then there exists a non-trivial solution for $\mu_{s}$, implying the existence of a giant component. In Figure 9, we show the region (marked in red) where a giant component emerges. On the other hand, the region (marked in blue) does not exist a giant component. As shown in Figure 9, it is observed that $R$ must be sufficiently large for both scenarios where $c$ is large and $c$ is small. Specifically, when $c$ is large (e.g., $c>2$), a large $R$ is necessary to surpass the percolation threshold for $p$ in (3). Surpassing this threshold leads to the learning of a large number of skills, thereby facilitating the automatic emergence of a giant component within the skill association graph due to a large number of skills in a text. Conversely, when $c$ is small (e.g., $c<1$), nearly every skill in a text is learned by the 1-skill learner (since the average number of skills per text is less than 1). This results in a very small $p$. Nonetheless, even in this scenario, a large $R$ is required to learn a sufficient number of skills to meet the threshold for a giant component’s existence as stipulated in (18). We also note from Figure 9 that the minimum $R$ to have a giant component is 1.1 and this is achieved when $c=1.29$.
V Hierarchy of skills
One common approach to train a domain-specific LLM is to adopt a pre-trained model, commonly referred to as a foundation model or a basic model, and fine tune it with additional domain-specific texts. Motivated by this, we consider the setting where there are two classes of skills: the class of basic skills ${\cal S}$ and the class of domain-specific skills ${\cal S}_{f}$ (see Figure 10 for an illustration). The original set of texts ${\cal T}$ is treated as a basic set of texts, which requires only basic skills to understand. In addition, we also consider an additional set of domain-specific texts ${\cal T}_{f}$, where each requires learning a certain number of domain-specific skills to understand. A prerequisite of learning a domain-specific skill requires learning a random number of basic skills first. Thus, there is a hierarchy of skills.
Figure 10: This figure is an illustration of the hierarchy of skills. The basic model is trained first and then fine-tuned with domain-specific texts. A prerequisite of learning a domain-specific skill requires learning a random number of basic skills.
In particular, following the procedure in Section III-A, we first train a basic model (foundation model) for a 1-skill learner using a set of basic texts ${\cal D}$ sampled from ${\cal T}$. Then, the probability that a randomly selected basic skill is learned can be computed as $\zeta=1-e^{-cR(1-p)}$, where $p$ is the solution to (3). Suppose that the basic skills required by a domain-specific skill are uniformly selected from the set of basic skills $𝒮\color[rgb]{0,0,0}{\cal S}$. Let $\delta_{\ell}$ be the probability that the number of basic skills required by a domain-specific skill is $\ell$ and $\Lambda_{f}(x)=\sum_{\ell=0}^{\infty}\delta_{\ell}x^{\ell}$ be the corresponding generating function. Then we say a domain-specific skill is learnable if all its basic skills are learned in the basic model. Hence, it occurs with probability
$\sum_{\ell=0}^{\infty}\delta_{\ell}{\zeta}^{\ell}=\Lambda_{f}(\zeta).$ (19)
To fine-tune the basic model, we collect a subset of domain-specific texts ${\cal D}_{f}$ and present these $|{\cal D}_{f}|$ texts to a 1-skill learner to acquire domain-specific skills. Let $R_{f}=|{\cal D}_{f}|/|S_{f}|$ be the ratio of the number of training domain-specific texts to the number of domain-specific skills. Similar to the basic model, we make the following two assumptions about sampled domain-specific texts:
(Poisson degree distribution) Denote by $N_{f}(t)$ the number of domain-specific skills in the domain-specific text $t\in{\cal D}_{f}$. Then $\{N_{f}(t),t\in{\cal D}_{f}\}$ are independent Poisson random variables with mean $c_{f}$, i.e., ${\bf\sf P}(N_{f}(t)=k)=e^{-c_{f}}\frac{c_{f}^{k}}{k!}$ for $k=0,1,2,\ldots$.
(Uniform connection) Each edge of a domain-specific text node is connected to a domain-specific skill node independently and uniformly. Thus, the probability that an edge of the domain-specific text node $t$ is connected to a particular domain-specific skill node $s$ is $1/|{\cal S}_{f}|$.
A domain-specific text node is said to be learnable if all of the domain-specific skill nodes connected to it are learnable. Otherwise, it is called a non-learnable domain-specific text node. We assume that nothing can be learned from a non-learnable domain-specific text when it is presented to a learner. From (A1d), (A2d) and (19), we deduce that the probability a randomly selected domain-specific text is learnable is given by
$\sum_{\ell=0}^{\infty}e^{-c_{f}}\frac{c_{f}^{\ell}}{\ell!}(\Lambda_{f}(\zeta))^{\ell}=e^{-c_{f}(1-\Lambda_{f}(\zeta))}.$ (20)
Define the bipartite graph formed by the learnable domain-specific skill nodes and the learnable domain-specific text nodes as the learnable domain-specific bipartite graph. In such a bipartite graph, there are approximately $\Lambda_{f}(\zeta)|{\cal S}_{f}|$ learnable domain-specific skill nodes and $e^{-c_{f}(1-\Lambda_{f}(\zeta))}|{\cal D}_{f}|$ learnable domain-specific text nodes. Furthermore, from Bayes’ formula, we deduce that the degree of a randomly selected learnable domain-specific text node $t$ in the learnable domain-specific bipartite graph follows a Poisson random variable with mean $c_{f}\Lambda_{f}(\zeta)$ since
[…]
$=\frac{e^{-c_{f}}\frac{c_{f}^{\ell}}{\ell!}(\Lambda_{f}(\zeta))^{\ell}}{e^{-c_{f}(1-\Lambda_{f}(\zeta))}}=e^{-c_{f}\Lambda_{f}(\zeta)}\frac{(c_{f}\Lambda_{f}(\zeta))^{\ell}}{\ell!}$ (21)
This leads to the following proposition:
Proposition 6
Consider the learnable domain-specific bipartite graph. Let $\tilde{\cal S}_{f}$ (resp. $\tilde{\cal D}_{f}$) be the set of learnable domain-specific skills (resp. texts). The learnable domain-specific bipartite graph satisfies the following properties:
(Poisson degree distribution) Denote by $\tilde{N}_{f}(t)$ the number of learnable domain-specific skills in the learnable domain-specific text $t\in\tilde{\cal D}_{f}$. Then $\{\tilde{N}_{f}(t),t\in\tilde{\cal D}_{f}\}$ are independent Poisson random variables with mean $c_{f}\Lambda_{f}(\zeta)$.
(Uniform connection) Each edge of a learnable domain-specific text node is connected to a learnable domain-specific skill node independently and uniformly. Thus, the probability that an edge of the learnable domain-specific text node $t$ is connected to a particular learnable domain-specific skill node $s$ is $1/|\tilde{\cal S}_{f}|$.
In light of Proposition 6, we can apply the same density evolution analysis used in Section III to analyze the learnable domain-specific bipartite graph. Suppose that we still use the 1-skill learner for fine tuning. Let $p_{f}$ be the probability that the text end of a randomly selected edge in the learnable domain-specific bipartite graph is not learned after fine tuning. Note that $|\tilde{\cal S}_{f}|\approx\Lambda_{f}(\zeta)|{\cal S}_{f}|$ and $|\tilde{\cal D}_{f}|\approx e^{-c_{f}(1-\Lambda_{f}(\zeta))}|{\cal D}_{f}|$. Replacing $c$ with $c_{f}\Lambda_{f}(\zeta)$ and $R$ with $R_{f}e^{-c_{f}(1-\Lambda_{f}(\zeta))}/\Lambda_{f}(\zeta)$ in (3) yields
$p_{f}=1-\exp\Big{(}{-c_{f}\Lambda_{f}(\zeta)e^{-(1-p_{f})c_{f}R_{f}e^{-c_{f}(1-\Lambda_{f}(\zeta))}}}\Big{)}.$ (22)
Similarly, let $\zeta_{f}$ be the probability that a randomly selected domain-specific skill node in the learnable domain-specific bipartite graph is learned after fine tuning. Then (cf. (4))
$\zeta_{f}=1-e^{-(1-p_{f})c_{f}R_{f}e^{-c_{f}(1-\Lambda_{f}(\zeta))}}.$ (23)
Now we derive the probability of the testing error of a randomly selected text from the domain-specific texts. Denote such a probability as $\epsilon_{f}$. Assume that this randomly selected domain-specific text has the same distribution as the training data, i.e., it satisfies (A1d) and (A2d). The testing error consists of two cases: (i) a randomly selected text from the set of domain-specific texts ${\cal T}_{f}$ is not learnable, and (ii) a randomly selected text from the set of domain-specific texts ${\cal T}_{f}$ is learnable but one of the domain-specific skills in that text is not learned. As shown in (20), the probability that a randomly selected text from the domain-specific texts is not learnable is $1-e^{-c_{f}(1-\Lambda_{f}(\zeta))}$. Since the degree distribution of a domain-specific text in the learnable domain-specific bipartite graph is Poisson with mean $c_{f}\Lambda_{f}(\zeta)$ in (21), the probability that one of the domain-specific skills in a learnable domain-specific text is not learned is (cf. (5))
$1-\sum^{\infty}_{k=0}e^{-c_{f}\Lambda_{f}(\zeta)}\frac{(c_{f}\Lambda_{f}(\zeta))^{k}}{k!}(\zeta_{f})^{k}$
$=1-e^{-c_{f}\Lambda_{f}(\zeta)(1-\zeta_{f})}.$ (24)
Thus, the probability for Case (ii) is
$(1-e^{-c_{f}\Lambda_{f}(\zeta)(1-\zeta_{f})})e^{-c_{f}(1-\Lambda_{f}(\zeta))}.$ (25)
Combining Cases (i) and (ii) in (V) and (25) leads to
$\epsilon_{f}$ $=$ $1-e^{-c_{f}(1-\Lambda_{f}(\zeta))}$ (26)
$+(1-e^{-c_{f}\Lambda_{f}(\zeta)(1-\zeta_{f})})e^{-c_{f}(1-\Lambda_{f}(\zeta))}.$
Figure 11: The contour plot of the probability of the testing error $\epsilon_{f}$ for the 1-skill learner as a function of $R$ and $R_{f}$ with values $c=3$ and $c_{f}=3$.
In view of (26), there are two thresholds for achieving low testing errors in a domain-specific task: (i) a large number of basic skills are learned when the number of basic texts exceeds the threshold in the foundation model, and (ii) a large number of domain-specific skills are learned when the number of domain-specific texts exceeds the threshold in the fine-tuning model. In Figure 11, we present a contour plot of the testing error $\epsilon_{f}$ as a function of $R$ and $R_{f}$ for the 1-skill learner with values of $c=3$ and $c_{f}=3$. The number of basic skills required by a domain-specific skill is Poisson distributed with mean $3$, in which case $\Lambda_{f}(x)=e^{-3(1-x)}$. The graph features two percolation thresholds associated with $R$ and $R_{f}$. It is observed that if $R<1$, the majority of domain-specific text nodes remain unlearned, primarily because the basic skills are not acquired, indicating the first percolation threshold. Conversely, when $R>1$, attention shifts to the second percolation threshold, $R_{f}$; in this scenario, a significant number of domain-specific texts are learned, provided that $R_{f}>1$. This distinction underscores the critical values of $R$ and $R_{f}$ in determining the learning outcomes within the graph.
VI Extensions to multiple classes of skills and texts
In this section, we extend the analysis to the setting with multiple classes of skills and texts, which is motivated by the existence of multiple subjects in texts, such as math, physics, chemistry, law, etc. Note that, in these settings, a text may require skills from multiple classes to understand and, thus, introduces interdependence among classes. However, different from Section V, the multiple classes of skills are not assumed to be hierarchical in nature and, thus, could be learned without a prerequisite on the knowledge of other classes of skills.
VI-A Poisson learners with multiple classes of skills/texts
We first extend the density evolution analysis for Poisson learners to settings with multiple classes of skills and texts. Our analysis is similar to the framework of Poisson receivers in [26].
VI-A1 The ensemble of random bipartite graphs
[…]
We make the following two assumptions about the sampled texts:
(Poisson degree distribution) Denote by $N_{k,j}(t)$ the number of class $k$ skills in a text $t$ in the set of class $j$ texts ${\cal D}_{j}$. Then $\{N_{k,j}(t),t\in{\cal D}_{j}\}$, $j=1,2,\ldots,J$, $k=1,2,\ldots,K$, are independent Poisson random variables with means $c_{k,j}$.
(Uniform connection) Each class $(k,j)$-edge of a text node $t\in{\cal D}_{j}$ is connected to a skill node in ${\cal S}_{k}$ independently and uniformly. Thus, the probability that a class $(k,j)$-edge from a text $t\in{\cal D}_{j}$ is connected to a particular skill node $s\in{\cal S}_{k}$ is $1/|{\cal S}_{k}|$.
[…]
Proposition 7
[…]
(ii) (Uniform connection) Each class $(k,j)$-edge of a class $k$ skill node is connected to a class $j$ text node independently and uniformly. Thus, the probability that a class $(k,j)$-edge from a skill node $s\in{\cal S}_{k}$ is connected to a particular class $j$ text node is $1/(\alpha_{j}|{\cal D}|)$.
To gain insight into such an extension, we note that the collection of class $(k,j)$-edges forms a bipartite graph with a single class of skills and a single class of texts. As such, the ensemble of bipartite graphs generated with $K$ classes of skills and $J$ classes of texts can be viewed as the union of $K\times J$ independent ensembles of bipartite graphs, each with a single class of skills and a single class of texts. In Figure 12, we show an illustration for a skill-text bipartite graph with two classes of skills and three classes of texts.
Figure 12: An illustration for a skill-text bipartite graph with two classes of skills and three classes of texts.
An edge is a class $k$ edge if the skill end of the edge is connected to a class $k$ skill. From Proposition 7, the probability that a class $k$ edge is a class $(k,j)$-edge is
$r_{k,j}=\frac{d_{k,j}}{d_{k}}=\frac{c_{k,j}\alpha_{j}}{\sum_{j^{\prime}=1}^{J}c_{k,j^{\prime}}\alpha_{j^{\prime}}}.$ (27)
VI-A2 Density evolution for Poisson learners with multiple classes of skills
We say a Poisson learner with $K$ classes of skills is subject to a Poisson offered load $\rho=(\rho_{1},\rho_{2},\ldots,\rho_{K})$ from a text $t$ if the number of class $k$ skills connected to the text node $t$, $k=1,2,\ldots,K$, are independent Poisson random variables with means $\rho_{k}$, $k=1,2,\ldots,K$. Analogous to the definition of a Poisson receiver with multiple class of input traffic in [27], we can define a Poisson learner with multiple class of skills as follows.
Definition 8
(Poisson learners with multiple classes of skills) An abstract learner is called a $(P_{{\rm suc},1}(\rho),P_{{\rm suc},2}(\rho),\ldots,P_{{\rm suc},K}(\rho))$-Poisson learner with $K$ classes of skills if the learner is subject to a Poisson offered load $\rho=(\rho_{1},\rho_{2},\ldots,\rho_{K})$ from a text, a tagged (randomly selected) class $k$ skill is learned from the text with probability $P_{{\rm suc},k}(\rho)$, for $k=1,2,\ldots,K$.
For a $(P_{{\rm suc},1}(\rho),P_{{\rm suc},2}(\rho),\ldots,P_{{\rm suc},K}(\rho))$-Poisson learner, the throughput of class $k$ skills (defined as the expected number of class $k$ skills that are learned) from a text with a Poisson offered load $\rho$ is thus given by
$\Theta_{k}=\rho_{k}\cdot P_{{\rm suc},k}(\rho),$ (28)
$k=1,2,\ldots,K$.
Suppose that $J$ classes of texts (a total of $|{\cal D}|$ texts) are presented to a Poisson learner with $K$ classes of skills. Our study utilizes the tree evaluation approach, as detailed in [23, 17, 29, 19], along with the concepts of reduced Poisson offered load, referenced in [27]. The tree evaluation method is also recognized as the density evolution analysis within the field of information theory [14, 24, 25]. Similar to the analysis in the original IRSA [17] and the coded Poisson receivers [26], we define $\Lambda_{k,\ell}$ as the probability that a class $k$ skill node has $\ell$ edges. From Proposition 7, we know that
$\Lambda_{k,\ell}=e^{-d_{k}}\frac{d_{k}^{\ell}}{\ell!},\;\ell=1,2,\dots$ (29)
Let
$\Lambda_{k}(x)=\sum_{\ell=0}^{\infty}\Lambda_{k,\ell}\cdot x^{\ell}$ (30)
be the generating function of the degree distribution of a class $k$ skill node, and
$\lambda_{k}(x)=\frac{\Lambda_{k}^{\prime}(x)}{\Lambda_{k}^{\prime}(1)}$ (31)
be the generating function of the excess degree distribution of a class $k$ skill node. Since the excess degree distribution of a Poisson degree distribution is also Poisson with the same mean (as it can be easily shown by using (31)), we have
$\Lambda_{k}(x)=\lambda_{k}(x)=e^{-d_{k}(1-x)}.$ (32)
Our density evolution analysis involves the following steps:
(i) Let $q_{k}^{(i)}$ be the probability that the skill end of a randomly selected class $k$ edge has not been learned after the $i^{th}$ SCNS iteration. The offered load of a class $j$ text to the Poisson learner after the $i^{th}$ SCNS iteration has a Poisson distribution with mean $q_{k}^{(i)}c_{k,j}$. As pointed out in [27], this is due to two important closure properties of Poisson random variables: (a) randomly splitting a Poisson random process yields independent Poisson random processes, and (b) the excess degree distribution of a Poisson random variable is Poisson with the same mean. Consequently, the Poisson offered load is reduced from $c_{k,j}$ to $q_{k}^{(i)}c_{k,j}$ after the $i^{th}$ SCNS iteration. Let
$\tilde{c}_{j}=(c_{1,j},c_{2,j},\ldots,c_{K,j}),$ (33)
and
$q^{(i)}=(q_{1}^{(i)},q_{2}^{(i)},\ldots,q_{K}^{(i)}).$ (34)
We can represent the offered load at a class $j$ Poisson receiver after the $i^{th}$ SCNS iteration by the vector $q^{(i)}\circ\tilde{c}_{j}$, where $\circ$ denotes the element-wise multiplication of two vectors. This reduced load argument is also the key step that greatly reduces the computational complexity in the density evolution method.
(ii) Let $p_{k,j}^{(i+1)}$ be the probability that the text end of a randomly selected class $(k,j)$-edge has not been learned after the $(i+1)^{th}$ SCNS iteration. Then
$p_{k,j}^{(i+1)}=1-P_{{\rm suc},k}(q^{(i)}\circ\tilde{c}_{j}).$ (35)
The above equation follows directly from the definition of a Poisson learner in Definition 8 as the offered load of a class $j$ text to the Poisson learner after the $i^{th}$ SCNS iteration is $q^{(i)}\circ\tilde{c}_{j}$.
(iii) Let $p_{k}^{(i+1)}$ be the probability that the text end of a randomly selected class $k$ edge has not been learned after the $(i+1)^{th}$ SCNS iteration. Since a class $k$ edge is a class $(k,j)$-edge with probability $r_{k,j}$, it follows that
$p_{k}^{(i+1)}$ $=$ $\sum_{j=1}^{J}r_{k,j}p_{k,j}^{(i+1)}$ (36)
$=$ $1-\sum_{j=1}^{J}r_{k,j}P_{{\rm suc},k}(q^{(i)}\circ\tilde{c}_{j}).$
(iv) The probability $q_{k}^{(i)}$ can be computed recursively from the following equation:
$q_{k}^{(i+1)}=\lambda_{k}\Big{(}1-\sum_{j=1}^{J}r_{k,j}P_{{\rm suc},k}(q^{(i)}\circ\tilde{c}_{j})\Big{)},$ (37)
with $q_{k}^{(0)}=1$. To see this, note that a skill end $s$ of a randomly selected edge is not learned after the $(i+1)^{th}$ iteration if the text nodes connected to $s$ (excluding the text end of the randomly selected edge) were not able to learn the skill $s$ after the $i^{th}$ iteration. Let $\lambda_{k,\ell}$ be the probability that the skill end of a randomly selected class $k$ edge has additional $\ell$ edges. Thus, the probability that the skill end of a randomly selected class $k$ edge cannot be learned after the $(i+1)^{th}$ iteration is
$q_{k}^{(i+1)}$ $=$ $1-\sum_{\ell=0}^{\infty}\lambda_{k,\ell}\cdot\Big{(}1-(p_{k}^{(i+1)})^{\ell}\Big{)}$ (38)
$=$ $\lambda_{k}(p_{k}^{(i+1)}).$
Using (36) in (38) yields (37). Note from (32) and (38) that
$q_{k}^{(i+1)}=e^{-d_{k}(1-p_{k}^{(i+1)})}.$ (39)
(v) Let $\tilde{P}_{{\rm suc},k}^{(i)}$ be the probability that a randomly selected class $k$ skill can be learned after the $i^{th}$ iteration. This is the probability that at least one text end of the edges connected to $s$ has been learned after the $i^{th}$ iteration. Since the probability that a randomly selected class $k$ skill node has $\ell$ edges is $\Lambda_{k,\ell}$, we have from (36) that
$\tilde{P}_{{\rm suc},k}^{(i)}=\sum_{\ell=0}^{\infty}\Lambda_{k,\ell}\cdot\Big{(}1-(p_{k}^{(i)})^{\ell}\Big{)}$
$=\sum_{\ell=0}^{\infty}\Lambda_{k,\ell}\cdot\Big{(}1-(1-\sum_{j=1}^{J}r_{k,j}P_{{\rm suc},k}(q^{(i-1)}\circ\tilde{c}_{j}))^{\ell}\Big{)}$
$=1-\Lambda_{k}\Big{(}1-\sum_{j=1}^{J}r_{k,j}P_{{\rm suc},k}(q^{(i-1)}\circ\tilde{c}_{j})\Big{)}.$ (40)
Here, we only outline the steps to derive the recursive equations. A rigorous proof for the density evolution analysis requires the concentration theorem (Theorem 2 of [25]) to prove the ergodicity of the system. i.e., as $|{\cal D}|\to\infty$, the average fraction of skill nodes that have not been learned after the $i^{th}$ SCNS iteration converges to the probability that a randomly selected skill node has not been learned after the $i^{th}$ SCNS iteration.
One can also represent the recursive equations using $p_{k}^{(i)}$’s. From (36) and (39), we also have for $k=1,2,\ldots,K$,
$p_{k}^{(i+1)}=1-\sum_{j=1}^{J}r_{k,j}P_{{\rm suc},k}(\gamma_{1}^{(i)},\gamma_{2}^{(i)},\ldots,\gamma_{K}^{(i)}).$ (41)
where
$\gamma_{k}^{(i)}=e^{-d_{k}(1-p_{k}^{(i)})}c_{k,j},$ (42)
$k=1,2,\ldots,K$. These $K$ coupled equations can be used for obtaining the scaling law for $p$ with respect to $R$, given the parameters $c_{k,j}$’s, $\alpha_{j}$’s, and $\beta_{k}$’s. Note that (41) reduces to (8) when $J=K=1$ and $i\to\infty$.
VI-B Deterministic $\psi$-learners
The framework of Poisson learners operates within a probabilistic context where skills are learned according to certain probabilities. As in [28], we can also define deterministic learners. Denote by ${\cal Z}^{+}$ the set of nonnegative integers. We say a learner with $K$ classes of skills is subject to a deterministic load $n=(n_{1},n_{2},\ldots,n_{K})\in{{\cal Z}^{+}}^{K}$ from a text if the number of class $k$ skills is $n_{k}$.
Definition 9
($\psi$-learner with multiple classes of skills [28]) Consider a deterministic function
$\psi:{{\cal Z}^{+}}^{K}\rightarrow{{\cal Z}^{+}}^{K}$
that maps a $K$-vector $n=(n_{1},n_{2},\ldots,n_{K})$ to the $K$-vector $(\psi_{1}(n),\psi_{2}(n),\ldots,\psi_{K}(n))$. An abstract learner is called a $\psi$-learner (with $K$ classes of skills) if the number of class $k$ skills that are learned is exactly $\psi_{k}(n)$, $k=1,2,\ldots,K$, when the learner is subject to a deterministic load $n=(n_{1},n_{2},\ldots,n_{K})$.
One typical example of the $\psi$-learner is 1-skill learner. As before, one can also generalize 1-skill learner to $|{\cal D}|$-skill learner, where there are at most $|{\cal D}|$ novel skills can be learned. The $|{\cal D}|$-skill learner is a $\psi$-learner with
$\psi(n)=\left\{\begin{array}[]{cc}n&\mbox{if}\;n\leq D\\ 0&\mbox{otherwise}\end{array}\right..$ (43)
Analogous to the proof of Theorem 14 in [28], it is shown that for every $\psi$-learner, there is an induced Poisson learner. This relationship is established by computing the throughput of the $\psi$-learner when it is subject to a Poisson offered load $\rho$, and then using (28) to determine the success probability function. For the $|{\cal D}|$-skill learner, the throughput for a Poisson offered load $\rho$ is given by
$\sum_{t=1}^{D}t\frac{e^{-\rho}\rho^{t}}{t!}=\rho\sum_{t=0}^{D-1}\frac{e^{-\rho}\rho^{t}}{t!}.$
From (28), it is a Poisson learner with the following success probability function:
$P_{\rm suc}(\rho)=\sum_{t=0}^{D-1}\frac{e^{-\rho}\rho^{t}}{t!}.$ (44)
To motivate the extension to multiple classes of skills, consider a learner with two classes of skills: class 1 and class 2. Similar to the 1-skill learner, this learner can learn a novel skill, whether class 1 or class 2, if it is the only novel skill in a text. Additionally, when a text contains both a class 1 novel skill and a class 2 novel skill, the learner can first learn the class 1 skill and then apply the SCNS strategy to reduce the number of novel skills to a single class 2 skill. Consequently, since the learner is capable of learning a single class 2 novel skill, it can also learn the class 2 skill in the text. Let $n_{1}$ and $n_{2}$ be the number of class 1 and class 2 skills, respectively, in a text. Then, such a deterministic learner is a $\psi$-learner with two classes of skills, where
$\psi(n_{1},n_{2})=\left\{\begin{array}[]{cc}(n_{1},n_{2})&\mbox{if}\;(n_{1},n_{2})\leq(1,1)\\ (0,0)&\mbox{otherwise}\end{array}\right..$ (45)
Since such a learning process is analogous to the near-far successive interference cancellation decoding scheme in [30], we refer to such a $\psi$-learner as a near-far learner. Now we can use the throughput formula in (28) to show that a near-far learner is also a Poisson learner with the following two success probability functions:
$P_{{\rm suc},1}(\rho_{1},\rho_{2})$ $=$ $e^{-\rho_{1}}(e^{-\rho_{2}}+\rho_{2}e^{-\rho_{2}}),$
$P_{{\rm suc},2}(\rho_{1},\rho_{2})$ $=$ $e^{-\rho_{2}}(e^{-\rho_{1}}+\rho_{1}e^{-\rho_{1}}).$ (46)
Suppose that there is only one class of texts, i.e., $J=1$. Now consider training with the near-far learner. Then we have from (41) that
$p_{1}^{(i+1)}=1-P_{{\rm suc},1}(e^{-d_{1}(1-p_{1}^{(i)})}c_{1},e^{-d_{2}(1-p_{2}^{(i)})}c_{2}),$
$p_{2}^{(i+1)}=1-P_{{\rm suc},2}(e^{-d_{1}(1-p_{1}^{(i)})}c_{1},e^{-d_{2}(1-p_{2}^{(i)})}c_{2}),$
where $d_{k}=(c_{1}+c_{2})R/\beta_{k}$, $k=1$ and 2.
Figure 13: The plot illustrates the computed values of $p_{1}$ (and $p_{2}$) for the near-far learner as a function of $R$, which ranges from 0.1 to 10 with a step size of 0.05, for each $\beta$ value (0.1, 0.3, 0.5, 0.7, 0.9). The average degrees of the two classes of skills are $c_{1}=3$ and $c_{2}=7$, respectively. The computation is based on the two recursive equations in (VI-B) with the initial condition $p_{1}^{(0)}=p_{2}^{(0)}=1$, where $P_{{\rm suc},k}(\rho)$, $k=1$ and 2, are defined in (VI-B). The number of iterations is 100.
In Figure 13, we plot the computed values of $p_{1}$ as a function of $R$ for each $\beta$ value (0.1, 0.3, 0.5, 0.7, 0.9). The average degrees of the two classes of skills are $c_{1}=3$ and $c_{2}=7$, respectively. For this example, it is noted that the values of $p_{2}$ overlap with those of $p_{1}$, making $p_{2}$ indistinguishable in the figure. Interestingly, thresholds are also observed for these five curves. As demonstrated in [26], the thresholds for these two classes of skills coincide due to the coupling of the density evolution equations (similar to those in (VI-B)), leading to the simultaneous emergence of various classes of skills.
VII Application to Semantic Compression and Communication
The lossless compression of a text requires exact recovery of the sequence of tokens in a text. However, if our interest lies in recovering only the semantic meaning of a text, we might be able to compress it using fewer bits than required for lossless compression. This concept is referred to as the semantic (level B) problem in [31].
Definition 10
(Semantic Compression) A compression method is termed a semantic compression method if the recovered text is semantically equivalent to the original text. An abstract learner is called generative if it can generate a text of tokens given a set of learned skills.
In this section, we assume that the abstract learners discussed in the previous section are also generative. Intuitively, to render an abstract learner generative, one can store a “table” of tokens corresponding to learned skills. Given a set of learned skills, the generative learner then uses the stored tokens to compose a text. It is known from [12] that transformers [11] are, in fact, Hopfield networks, which are a form of associative memory. In other words, transformers are capable of storing tables of tokens for learned skills.
[…]
For semantic communication, we can utilize a semantic encoder/decoder. In Figure 14, we present the general architecture of a semantic communication system, which consists of a semantic transmitter, a semantic receiver, and a physical channel. The semantic transmitter includes a semantic encoder and a channel encoder, while the semantic receiver includes a semantic decoder and a channel decoder. The communication between the channel encoder and decoder occurs over the physical channel, referred to as level A communication in [31]. Conversely, the communication between the semantic encoder and decoder is termed level B communication in [31].
Recent literature on semantic communications, such as [33, 34, 35], has proposed an end-to-end approach for jointly training the semantic and channel encoder/decoder. This method is claimed to be superior to separate training. However, the end-to-end approach does not scale efficiently with data size. In fact, as shown in the previous section, a large dataset is necessary for training to exhibit the emergence of semantic capability. Hence, for the transmission of texts from a general semantic language, an LLM model is required at both the semantic transmitter and receiver, which would be difficult to retrain and adapt to varying physical channels. In light of this, a modular design (i.e., separate source and channel coding) may be more effective for semantic communication in practice.
Figure 14: The general framework of a semantic communication system.
For domain-specific texts discussed in Section V, semantic compression can also be used once fine-tuning is complete. If the number of domain-specific skills is much smaller than the number of basic skills, i.e., |$𝒮|\color[rgb]{0,0,0}{\cal S}_{f}|\ll|{\cal S}|$, one expects the compression gain for domain-specific texts in the fine-tuning model will surpass that achieved using the foundation model.
VIII Conclusions and Future Directions
Inspired by [13], we defined a semantic language as a skill-text bipartite graph. Skills can be learned by repeatedly presenting training texts to abstract learners, such as 1-skill learners and Poisson learners. By using density evolution analysis, we showed the emergence of learned skills when the ratio $R$ exceeds a certain threshold. Moreover, we derived the corresponding scaling law for the testing error. We also demonstrated how trained abstract learners can be used for semantic compression and semantic communication.
The present work is intended to inspire further thoughts on the emergence of learned skills in today’s LLMs, drawing relation to the iterative decoding procedure in communications and information theory. However, we recognize that there are still significant gaps between our theory and real-world systems. Several key factors that should be further investigated include:
  • (i)
    Modeling abstract learners: Real-world LLMs, such as GPT-4 [1] and Gemini [2], are transformer-based and are trained using the gradient descent algorithm. It is not clear how a skill is learned when a text of tokens is fed into these transformers.
  • (ii)
    Order of skills: If we view a skill as a thought, often a text can be represented by a chain (sequence) of thoughts. As such, the order of the skills appearing in a text could also be an important factor, thus, learning the mapping $\phi$ alone might not suffice for effective semantic encoding/decoding.
  • (iii)
    Soft learning: In our model for an abstract learner, a skill is learned in a definite manner. Exploring a soft learning setting, where a skill is partially learned, could be beneficial. For transformer-based LLMs, a softmax function is used for the next token prediction. Extending our theory to encompass soft learning might help bridge the gap between theory and real-world systems.
  • (iv)
    Scaling laws of the number of parameters: The scaling law obtained in this paper only pertains to the size of training texts. It is unclear how the testing error scale with the number of parameters. For instance, GPT-3 is known to have more than 1,750 billion parameters, and LLMs prior to GPT-3 do not exhibit the emergence of skills. One possible explanation for the necessity of a larger number of parameters is that the capacity of a transformer, defined as the number of skills (or patterns) that can be stored in the corresponding Hopfield network [12], depends on the number of parameters in the transformer.
  • (v)
    Experimental verification: Performing experiments to verify our theory in the context of real-world semantic languages is difficult. However, it may be possible to construct a semantic language, e.g., a simple programming language, with a small number of skills and employ transformer-based LLMs with a relatively small number of parameters to verify the theory.
Appendix A
In Table I, we provide a list of notations used in this paper.
TABLE I: List of Notations
${\cal A}$ A set of tokens (symbols)
$c$ The average number of skills in a (basic) text (see (A1))
$c_{f}$ The average number of skills in a domain-specific text (see (A1d))
$c_{k,j}$ The average number of class $k$ skills in a class $j$ text (see (A1m))
${\tilde{c}}_{j}$ The vector of $c_{k,j}$, ${\tilde{c}}_{j}=(c_{1,j},c_{2,j},\ldots,c_{K,j})$
${\cal D}$ The set of training texts
${\cal D}_{f}$ The set of domain-specific texts for training
${\cal D}_{j}$ The set of class $j$ training texts
$\tilde{\cal D}_{f}$ The set of learnable domain-specific texts
$d_{k,j}$ The average number of class $j$ texts connected to a class $k$ skill node
$d_{k}$ The mean degree of a class $k$ skill node
$E$ The edge set of the skill-text bipartite graph
$G$ The skill-text bipartite graph
$i$ The number of SCNS iteration
$J$ The number of classes of texts
$j$ The index for the $j^{th}$ class of texts
$K$ The number of classes of skills
$k$ The index for the $k^{th}$ class of skills
${\cal L}$ A semantic language
$M(s)$ The number of edges connected to a skill node $s$
$M_{k,j}(s)$ The number of class $(k,j)$-edges connected to a class $k$ skill node
$N(t)$ The number of edges connected to a text node $t$
$N_{f}(t)$ The number of edges connected to a domain-specific text node $t$
$\tilde{N}_{f}(t)$ The number of edges connected to a learnable domain-specific text node $t$
$N_{k,j}(t)$ The number of class $(k,j)$-edges connected to a class $j$ text node
$p_{f}$ The probability that the text end of a randomly selected edge in the learnable domain-specific bipartite graph is not learned after fine tuning
$p_{G}$ The probability that a randomly selected skill node is in the giant component
$p^{(i)}$ The probability that the text end of a randomly selected edge is not learned after the $i^{th}$ iteration
$p_{k}^{(i)}$ The probability that the text end of a randomly selected class $k$ edge has not been learned after the $i^{th}$ SCNS iteration
$p_{k,j}^{(i)}$ The probability that the text end of a randomly selected class $(k,j)$-edge has not been learned after the $i^{th}$ SCNS iteration
$P_{\rm suc}(\rho)$ The success probability of a Poisson learner when the number of skills in a text is Poisson distributed with mean $\rho$
$P_{\text{suc},k}(\rho)$ The probability that a class $k$ skill is learned when the number of skills in a text is subject to a Poisson offered load $\rho$
$\tilde{P}_{{\rm suc},k}^{(i)}$ The probability that a randomly selected class $k$ skill can be learned after the $i^{th}$ SCNS iteration
$q^{(i)}$ The probability that the skill end of a randomly selected edge is not learned after the $i^{th}$ SCNS iteration
$q_{k}^{(i)}$ The probability that the skill end of a randomly selected class $k$ edge has not been learned after the $i^{th}$ SCNS iteration
$R$ The ratio of the number of training texts to the number of skills
$R_{f}$ The ratio of the number of training domain-specific texts to the number of domain-specific skills
$r_{k,j}$ The probability that a class $k$ edge is a class $(k,j)$-edge
${\cal S}$ The set of skills
${\cal S}_{k}$ The set of class $k$ skills
${\cal S}_{f}$ The set of domain-specific skills
$\tilde{\cal S}_{f}$ The set of learnable domain-specific skills
${\cal T}$ The set of texts composed of sequences of tokens
${\cal T}_{j}$ The set of class $j$ texts
${\cal T}_{f}$ The set of domain-specific texts
$\tilde{\cal T}_{f}$ The set of learnable domain-specific texts
${\cal Z}^{+}$ The set of nonnegative integers
$\alpha_{j}$ The ratio of the number of class $j$ training texts to the total number of training texts
$\beta_{j}$ The ratio of the number of class $k$ skills to the total number of skills
$\epsilon$ The probability of testing error of a randomly selected text
$\epsilon_{f}$ The probability of testing error of a randomly selected domain-specific text
$\Theta_{k}$ The throughput of class $k$ skills from a text with a Poisson offered load $\rho$
$\Lambda_{f}$ The generating function of the degree distribution of a domain-specific skill node
$\Lambda_{k,\ell}$ The probability that a class $k$ skill node has $\ell$ edges
$\Lambda_{k}(x)$ The generating function of the degree distribution of a class $k$ skill node
$\Lambda_{k}^{\prime}(x)$ The derivative of $\Lambda_{k}(x)$
$\Lambda_{k}^{\prime}(1)$ The mean degree of a class $k$ skill node
$\lambda_{k}(x)$ The generating function of the excess degree distribution of a class $k$ skill node
$\mu_{s}$ The probability that a skill node is connected to a small component via one of its edges
$\mu_{t}$ The probability that a text node is connected to a small component via one of its edges
$\rho$ The Poisson offered load $\rho=(\rho_{1},\ldots,\rho_{K})$
$\rho_{\ell}$ The probability of the number of basic skills required by a domain-specific skill is $\ell$
$\phi$ A function that maps a text $t$ in $T$ to a set of skills
$\psi$ A deterministic function that maps a $K$-vector $n=(n_{1},n_{2},\ldots,n_{K})$ to the $K$-vector $(\psi_{1}(n),\psi_{2}(n),\ldots,\psi_{K}(n))$
$\zeta$ The probability that a randomly selected skill is learned
$\zeta_{f}$ The probability that a randomly selected domain-specific skill is learned